OpenStudy (anonymous):

1/4c greater than or equal to 4 explain.

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }c = 16 > 4 \]

OpenStudy (anonymous):

well it has a line under the sign i just cant get it to show

OpenStudy (anonymous):

\[\le or \ge \]

OpenStudy (anonymous):

Like that ?

OpenStudy (anonymous):

yeah the second one

OpenStudy (anonymous):

That's what I was saying. The line under the greater than sign means that it's a true statement. So as I was saying \[\frac{ 1 }{ 4 }c = 16 > 4 ~ or ~you~could~say ~\frac{ 1 }{ 4 }c = 16 \ge4 \]

OpenStudy (anonymous):

oh yeah the second one

OpenStudy (anonymous):

The only difference is that the line means that it's a true statement. If the equation in the problem used the greater than symbol with the line under it, I'd say to go with . . . \[\frac{ 1 }{ 4 }c = 16 \ge 4 \]

OpenStudy (anonymous):

so i put that? ^

OpenStudy (anonymous):

Is the question multiple choice or written anwser?

OpenStudy (anonymous):

written answer

OpenStudy (anonymous):

Then yes, that would be the anwser.

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

My Pleasure!